Please help: The length of a rectangle is 4 feet less than twice the width of the rectangle. If the perimeter of the rectangle is 136 feet, find the length of the rectangle.
Answer by checkley71(8403)

LENGTH=2W-4 SO THE PERIMETER IS
2W+2(2W-4)=136
2W+4W-8=136
6W=136=8
6W=144
W=144/6
W=24 SOLUTION
THEN THE LENGTH IS 2*24-4=48-4=44
PROOF
2*24+2*44=136
48+88=136
136=136